Blue water gas is obtained by passing steam over red hot carbon enriched
coal (coke). The gas is called "blue" water gas because of its blue flame, a characteristic of the combustion CO. A typical composition shows 4.7% CO2, 41% CO, 49% H2, 0.8% CH4, and 4.5% N2. If the blue water gas at 30 deg C, 700 torrs and 93% RH is burned in wet air, 38% in excess, supplied at the same condition as BWG with 85% RH, calculte
A.) Orsat analysis of the stack gas if the molar ratio of CO2 to CO is 1:8 and H2 to CO is 1:4
B.) GCV of BWG
C.) % Calorific value lost due to sensible heat
D.) Thermal Efficiency
1
Expert's answer
2021-07-14T05:48:31-0400

Part a


"T= 25^0 C = 298 K\\
P_s= e^{73.649- \frac{7258.2}{298}-7.3037 \ln(298)+4.1653*10-6(298^2))(\frac{760}{101325})}=23.57 torr\\
N \space \space Dry \space \space BWD =100(\frac{745-23.57*0.9}{760})(\frac{273}{298})=87.25 kgmol \space \space BWG\\
\frac{nH_2O}{ n \space dry \space gas}=\frac{ p H_2O}{ p \space dry \space gas}\\
nH_2O = 87.24 (\frac{ 23.57*0.9}{745-23.57*0.9})=2.556 kgmole\\
O_2 \space theo = 46.5 + \frac{101.2}{4}-25.2 = 46.6 kgmole\\
O_2 \space supplied = 1.30*46.6=60.58 kgmole\\
N_2 \space air =60.58 + \frac{79}{21} = 227.896 kgmole\\
N_2 \space total = N_2 \space air + N_2 \space fuel = 227.896 +4.5=232.4 kgmole\\
CO_2 \space formed = \frac{8}{9}*46.6 = 41.33 kgmole\\
CO \space formed = \frac{1}{9}*46.6 = 5.167 kgmole\\
Unburned \space H_2 = \frac{1}{4}*5.167 = 1.2967 kgmole\\
H_2 \space combusted = 101.2( \frac{1}{2})-1.29167 = 49.3083 kgmole=n H_2O \space combustion\\
NH_2O \space from \space air= (60.58+227.896)(\frac{23.57*0.85}{745-23.57*0.85} = 7.9721 kgmole\\
Free \space O_2 from = 13.98+\frac{5.167}{2}+\frac{1.29167}{2} = 17.2092 kgmole\\"


Part b

"M^3 \space air /m^3 \space BWG = 2.8848* \frac{22.4}{1}* \frac{760}{745-23.57*0.9}*\frac{298}{273}*\frac{1}{22.4}*\frac{273}{298}* \frac{745-23.57*0.9}{760}=2.88"


Part c

"\frac{m^3 \space stack \space gas }{100 \space moles \space BWG} = 357.23 * \frac{22.4}{1}*\frac{101.325}{100}* \frac{573}{273}=17017.84 \frac{m^3}{kg mol}"


Part d

"\frac{T_C}{T_H}*100\\
\frac{25}{30}*100\\
83.33\%"